Complete the following table of data for Na2HAsO4. (Assume the density of water is 1.00 g/mL.)

T (°C) Solubility
(g/100. g)
Concentration
(M)
Ksp ΔG
(kJ/mol)
42.0 49.0 ? ? ?
48.0 54.9 ? ? ?
54.0 60.8 ? ? ?
60.0 66.8 ? ? ?
66.0 72.7 ? ? ?
72.0 78.6 ? ? ?
78.0 84.5 ? ? ?

Use a graph of ΔG vs. T to determine the following. (Enter your answers to three significant figures.)

What is the value of ΔH for the system?(in kJ/mol)

What is the value of ΔS for the system? (in J/mol·K)
Answer #1

Na2HAsO4 dissociates as

Na2HAsO4 2 Na+ + HAsO4-

hence, Ksp = [Na+ ] 2* [ HAsO4-]

let molar solubility of Ksp = S (M)

then [Na+ ] = 2S (M ) and  [ HAsO4-] = s (M)

hence,

Ksp = (2S)2* S = 4s3 M3

Molar mass of Na2HAsO4 = 185.9 g/mol

then,

molarity = (mass of Na2HAsO4 * 1000) / (molar mass of Na2HAsO4 * volume of water )

density is 1.00 g/L , hence mass of water = volume of water.

molar concentration (M) , Ksp and   at different temperature are calculated below

420 c

M = 49*1000/185.90*100 = 2.647

Ksp = 4*(2.647)3 = 74.21 M3

= - RTl n Ksp = -8.314*(42+273) ln 74.21 = - 11279.38 J = -11.279 KJ/mol

480 c

M = 54.9*1000/185.90*100 = 2.966

Ksp = 4*(2.966)3 = 10.4.36 M3

= - RTln Ksp = -8.314*(48+273) ln 104.36 = -12404.144 J = -12.40 KJ/mole

540 c

M = 60.8*1000/185.90*100 = 3.284

ksp =4*(3.284)3 = 141.66

= - RTln Ksp = -8.314*(54+273) ln 104.36 = -12635.99 J = -12.63 KJ/mole

600 c

M = 66.8*1000/185.90*100 = 3.609

ksp = 4*(3.609)3 = 188.02 M3

= - RTlnKsp = -8.314*(60+273) ln 188.02 = - 14497.70 J = -14.49 KJ/mole

660 c

M = 72.7*1000/185.09*100 = 3.927

Ksp = 4*(3.927)3 = 242.23 M3

= - RTlnKsp = -8.314*(66+273) ln 242.23 = -15472.95 J = -15.472 KJ/mole

720 c

M = 78.6*1000/185.09*100 = 4.246

Ksp = 4*(4.246)3 = 306.19 M3

= - RTlnKsp = -8.314*(72+273) ln 306.19 = - 16418.911 J = -16.41 KJ/mole

780 c

M = 84.5*1000/185.09*100 = 4.56

Ksp = 4*(4.56)3 = 379.275 M3

= - RTlnKsp = - 8.314*(78+273) ln 379.275 = -17329.11 J = -17.32 KJ/mole

now plot of vs T

the bestline equation is

y = -172.7 x +43222

comparing with the equation

= - T

where. slope = = 172.7 J/mol

and = intercept = 43222 KJ/mole = 43.22 KJ/mole